Skip to main content

IT Status Alerts

All Closed Alerts

Go to Page: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70 71 72 73 74 75 76 77 78 79 80 81 82 83 84 85 86 87 88 89 90 91 92 93 94 95 96 97 98 99 100 101 102 103 104 105 106 107 108 109 110 111 112 113 114 115 116 117 118 119 120 121 122 123 124 125 126 127 128 129 130 131 132 133 134 135 136 137 138 139 140 141 142 143 144 145 146 147 148 149 150 151 152 153 154 155 156 157 158 159 160 161 162 163 164 165 166 167 168 169 170 171 172 173 174 175 176 177 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 193 194 195 196 197 198 199 200 201 202 203 204 205 206 207 208 209 210 211 212 213 214 215 216 217 218 219 220 221 222 223 224 225 226 227 228 229 230 231 232 233 234 235 236 237 238 239 240 241 242 243 244 245

4882 results found, displaying results 361 - 381.

Scheduled Service Change: PeopleSoft: Emergency Migration request for Audit Reporting 4/14/2017

Start:
2017-04-14 04:30:00
Status:
Closed
Brief Description:
PeopleSoft: Emergency Migration request for Audit Reporting 4/14/2017
more details

Performance Issue: Delays for Lyris E-list

Start:
2017-04-13 09:00:00
Status:
Closed
Brief Description:
There have been delays for e-list (Lyris) messages since about 9am this morning.
Current Status:
The e-list service is back in normal operation. As of 11am, all backlogged e-list email has been processed and sent out.
more details

Scheduled Service Change: PeopleSoft: Bi-Weekly Migration of online and batch objects 4/13/2017

Start:
2017-04-13 05:00:00
Status:
Closed
Brief Description:
PeopleSoft: Bi-weekly migration of online and batch objects.
more details

Unplanned Outage: Wired and Wi-Fi data network outage in some parts of Bethe House N

Start:
2017-04-13 06:56:00
Status:
Closed
Brief Description:
-
Current Status:
The bethehousen4-3500 network switch is now back up and network connectivity has been restored to the affected areas of Bethe House.
more details

Unplanned Outage: Wired and Wi-Fi data network outage in Vet Research Tower

Start:
2017-04-10 09:28:00
Status:
Closed
Brief Description:
CIT has received reports that Vet Research Tower switch (vrt3-5400) is down.
Current Status:
The Vet Research Tower switch (vrt3-5400 is down due to construction.
more details

Scheduled Service Change: Upgrade Shibboleth IDP to latest version

Start:
2017-04-09 07:00:00
Status:
Closed
Brief Description:
On Sunday April 9th from 7:00 am - 9:00 am, Shibboleth Identity Provider will be upgraded to the latest version. There should be no outage to the service.
more details

Scheduled Service Change: Provisioning changes in support of forthcoming email changes

Start:
2017-04-07 05:30:00
Status:
Closed
Brief Description:
Identity Mgmt will make a broad set of relatively minor changes to the IdM database, both the Enterprise Directory and Active Directory, WhoIAm, HelpHero, and the provisioning processes, then populate a new Directory attribute cornelleduprovisionaccts.
more details

Scheduled Service Change: Planned maintenance for Adobe Sign

Start:
2017-04-08 08:00:00
Status:
Closed
Brief Description:
Adobe has notified us of planned downtime for Adobe Sign. The service will be unavailable in the window indicated.
more details

Performance Issue: Skype for Business Connection Issue

Start:
2017-03-31 13:32:00
Status:
Closed
Brief Description:
There have been reports of issues connecting with Skype for Business, with a message that says the server is temporarily unavailable. If you are experiencing problems connecting with Skype for Business, please contact the IT Service Desk.
Current Status:
Recent incidents appear to be part of a larger known issue, the workaround for which is available at https://it.cornell.edu/skype-win/skype-windows-problem-server-temporarily-unavailable . If anyone encounters a similar problem and the workaround doesn't fix it, they are encouraged to contact the IT Service Desk.
more details

Performance Issue: Partial Connectivity Issues with the Production PeopleSoft Instance.

Start:
2017-04-03 08:00:00
Status:
Closed
Brief Description:
Dear PeopleSoft Community, We are experiencing partial network connectivity issues with the Production system.  Users can access the system directly, but some access such as jobs run by Tidal and direct connections to the database are unavailable.
Current Status:
The partial network connectivity issues with PeopleSoft has been resolved.
more details

Unplanned Outage: MyIT, SmartIT and Classic Remedy currently not available

Start:
2017-04-02 21:30:00
Status:
Closed
Brief Description:
Vendor currently investigating the issue.
Current Status:
All Remedy applications are now available. Vendor applied a patch that fix the issue.
more details

Scheduled Service Change: Hardware upgrade: cornellnet2

Start:
2017-04-02 06:00:00
Status:
Closed
Brief Description:
Replace existing cornellnet2 Cisco router with new cornellnet2 Juniper router.
more details

Scheduled Service Change: Maintenance on Bomgar Remote Assistance server

Start:
2017-04-01 07:00:00
Status:
Closed
Brief Description:
Maintenance on Bomgar appliance to configure failover settings
more details

Unplanned Outage: Wired and Wi-Fi data network outage in Morrill Hall

Start:
2017-03-31 13:59:00
Status:
Closed
Brief Description:
CIT has received reports that Morrill Hall switch (morrill1-5400) is down.
Current Status:
Wired and Wi-Fi data network has been restored to Morrill Hall. The switch (morrill1-5400) is back up.
more details

Unplanned Outage: Wired and Wi-Fi data network outage in some parts of BeckerHouse

Start:
2017-03-30 07:21:00
Status:
Closed
Brief Description:
Wired and Wi-Fi data network outage in some parts of BeckerHouse.
Current Status:
The switches that were down at Becker House are now back up and Wired and Wi-Fi data network services have been restored.
more details

Scheduled Service Change: PeopleSoft Updates to PSPROD

Start:
2017-03-30 04:30:00
Status:
Closed
Brief Description:
PeopleSoft: Bi-weekly migration of online and batch objects and implementing the PUM Image 4 updates during this maintenance window.
more details

Performance Issue: SharePoint On-Premise Workflows

Start:
2017-03-22 10:25:00
Status:
Closed
Brief Description:
Workflows were not functioning in Sharepoint On-Premise site collections. A potential fix has been put in place.
Current Status:
Workflows and SharePoint On Premise have been exhibiting stable performance since Saturday, when the fix was put in place.
more details

Performance Issue: Desktop Everywhere Service (VDI) issues only for non-Kiosk subscribers

Start:
2017-03-29 10:40:00
Status:
Closed
Brief Description:
The Desktop Everywhere Service (VDI) is experiencing a problem with non-Kiosk subscribers.  
Current Status:
The issue with the Desktop Everywhere (VDI) service has been resolved and has been returned to full working order.  Those affected will be receiving a more detailed notice soon.
more details

Scheduled Service Change: Remedy Application OS Maintenance

Start:
2017-03-26 12:00:00
Status:
Closed
Brief Description:
BMC will be applying the latest patches recommended by Microsoft on the Remedy OnDemand systems.
more details

Scheduled Service Change: Kronos: Planned Outage MS Monthly Updates - 3/25/2017

Start:
2017-03-25 07:00:00
Status:
Closed
Brief Description:
Kronos: planned outage MS monthly updates
more details
Go to Page: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70 71 72 73 74 75 76 77 78 79 80 81 82 83 84 85 86 87 88 89 90 91 92 93 94 95 96 97 98 99 100 101 102 103 104 105 106 107 108 109 110 111 112 113 114 115 116 117 118 119 120 121 122 123 124 125 126 127 128 129 130 131 132 133 134 135 136 137 138 139 140 141 142 143 144 145 146 147 148 149 150 151 152 153 154 155 156 157 158 159 160 161 162 163 164 165 166 167 168 169 170 171 172 173 174 175 176 177 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 193 194 195 196 197 198 199 200 201 202 203 204 205 206 207 208 209 210 211 212 213 214 215 216 217 218 219 220 221 222 223 224 225 226 227 228 229 230 231 232 233 234 235 236 237 238 239 240 241 242 243 244 245